OpenStudy (anonymous):

-45r^4+20r^3-25^2+4r+17/-5r perform the division

OpenStudy (anonymous):

\[-45r^4+20r^3-25r^2+4r+17/-5r\]

OpenStudy (anonymous):

that is the right problem i was missing the r on the 25r^2

OpenStudy (anonymous):

__________________________________ -5r ) \(-45r^4 + 20 r^3 - 25 r^2 + 4r +17\) The first step is to figure out how many times '-5r' goes into '\(-45r^4\)'. In other words, what is \[\frac{-45r^4}{-5r}\]?

OpenStudy (anonymous):

is it 9?

OpenStudy (anonymous):

That's part of it, yep. You're missing some r's though.

OpenStudy (anonymous):

oh sorry 9r^2

OpenStudy (anonymous):

I'd say it's 9r^3, but that's close enough I think! So, next you write down: \(9r^3\) _____________________________________ -5r ) \(-45r^4+20r^3-25r^2+4r+17\) \(+45r^4\) That +45r^4 is 9r^3 times -5r, then the sign gets reversed so you can add the two numbers together to get a remainder. In this case, -45r^4, and +45r^4 add together to get 0, so there's no remainder. Next, how many times does -5r go into 20r^3?

OpenStudy (anonymous):

\[-4r^3\]

OpenStudy (anonymous):

Close! The -4 is correct, but r^3 divided by r is r^2. Remember:\[\frac{r^n}{r^m}=r^{n-m}\]In this case \[\frac{r^3}{r^1}=r^{3-1}=r^2\]

OpenStudy (anonymous):

ok now whats next? is it -25r^2/-5r?

OpenStudy (anonymous):

So next we write down: \(9r^3 - 4r^2\) ______________________________________ -5r )\(-45r^4+20r^3-25r^2+4r+17\) \(+45r^4\) \(0 + 20r^3\) \(- 20r^3\) Again, there's no remainder. You're right about what's next. What's -25r^2 divided by -5r?

OpenStudy (anonymous):

is it 5r^2 or just 5r?

OpenStudy (anonymous):

It's 5r. Next, we have a weird one. What's 4r divided by -5r?

OpenStudy (anonymous):

hmmm is it 0.8?

OpenStudy (anonymous):

Correct! Since that's less than 1, we say that -5r 'goes into' 4r 0 times, with a remainder of 4r. That looks like: \(+0\) ___________ -5r ) ....\(4r + 17\) 0 So now you have to divide the whole expression '4r + 17' by -5r. Again, it doesn't 'go in' very well. So that means something weird happens. Your whole division must be expressed as: \[\frac{-45r^4+20r^3-25r^2+4r+17}{-5r}=9r^3-4r^2+5r + \frac{4r+17}{-5r}\]Yeah, it looks pretty weird, but them's the breaks!

OpenStudy (anonymous):

many thanks man.